given cost and price functions c(q)=120q+43,000 and p(q)= -1.8q+870, what price should be set to maximum profit?

it should be $____ per item
Answer #1

Given and C(q) = 1209 +43000 p(q) = -1.89 +070 . Now R(q) = 6(e).q =((-1.8)q+870) a = -1.89² +3709 : Pra)= R(q)- c(a) Profitor maximum protit = 35125 so of should be ? be $35125 es per item Thank